Circular No. 7095

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990123 Further to the item on IAUC 7094, M. Feroci and L. Piro, Istituto di

Astrofisica Spaziale, CNR, Rome; F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, and University of Ferrara; V. Torroni and M. Smith, BeppoSAX Science Operation Center, Telespazio, Rome; and J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, provide the following elaboration of the information given on GCN 199 and 202: “A γ-ray burst triggered the BeppoSAX Gamma-Ray Burst Monitor (GRBM) on Jan. 23.40780 UT. The burst was also detected in the Wide Field Camera (WFC) 1. The GRBM lightcurve shows a multipeak complex structure that lasts for ∼ 100 s with two major peaks. The intensity of the highest peak is 11 200 counts/s in the band 40–700 keV. A preliminary estimation of the fluence is (3.5 ± 0.4) × 10^{−4} erg cm^{−2} s^{−1}. This is the brightest burst detected so far by the BeppoSAX GRBM simultaneously with the WFC. The WFC lightcurve, which lasts for a comparable time, is also complex, with one peak followed by a structured, higher plateau. The WFC peak intensity corresponds to a flux of 3.4 Crab in the energy band 1.5–26 keV, and it is reached ∼ 40 s after the GRBM peak. The position of the x-ray counterpart to GRB 990123, as derived by the quicklook WFC image, is α = 15^{h}25^{m}29^{s}, δ = +44^{o}45′.5 (equinox 2000.0), with an error radius of 2′ (99-percent confidence level). A follow-up observation with BeppoSAX narrow-field instruments is in progress.”

J. Zhu and H. T. Zhang, on behalf of the Beijing Astronomical Observatory (BAO) GRB team, write: “R-band images of the BeppoSAX WFC error circle of GRB 990123 were obtained on Jan. 23.756 UT (8.5 hr after the GRB) with the BAO 0.6-m Schmidt telescope at Xinglong. Due to bad weather, only one 20-min exposure image taken in thin cloud was usable (FWHM = 6′′). The central part of the image is posted at

http://vega.bac.pku.edu.cn/~zj/grb/grb990123.gif. A faint object

could be seen close to the optical candidate position suggested by S. C. Odewahn et al. (IAUC 7094); its position from our measurement is α = 15^{h}25^{m}30^{s}.28, δ = +44^{o}45′59′′.0 (1σ = 0′′.5), with a magnitude of 19.2 (± 0.5?) from four stars in the USNO-A V1.0 catalogue. The object in our image seems slightly extended and slightly to the southwest, compared with the discovery image of Odewahn et al., but it seems to be impossible to confirm from only one image of poor quality. However, it could be concluded that if such an object (in our image) is not real, the optical transient in Odewahn et al.’s images must have faded greatly during the 4.6-hr interval.”

1999 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green